# D-tensor paraproducts and its caricatures

### v1

@@ 1 Introduction
$\leq Cd_{1}(x_{1},\tilde{x}_{1})^{2\alpha}d_{2}(x_{2},\tilde{x}_{2})^{2\alpha}\ldots d_{d}(x_{d},\tilde{x}_{d})^{2\alpha}$ (64)
since $\Delta(A,f)$ is piecewise constant on dyadic boxes (see 59). Now consider the case where $x,x^{\prime}$ are not contained in same dyadic box. Again, summing over the scaling and location parameters in $\mathcal{L}$ for $\Delta(A,f)$ one has:
@@
$\leq Cd_{1}(x_{1},x^{\prime}_{1})^{2\alpha}d_{1}(x_{2},x^{\prime}_{2})^{2\alpha}\ldots d_{d}(x_{d},x^{\prime}_{d})^{2\alpha}$ (65)
where we order the scales in $\mathcal{L}$ such that $\mathbf{j}^{(0)}\leq\mathbf{j}^{(1)}\leq\ldots\leq\mathbf{j}^{(N)}$ indicates $2^{-(\sum_{l}^{d}\mathbf{j}_{l}^{(0)})}\leq 2^{-(\sum_{l}^{d}\mathbf{j}_{l}^{(1)})}\leq\ldots\leq 2^{-(\sum_{l}^{d}\mathbf{j}_{l}^{(N)})}$. The first inequality holds from the estimates on $\Delta(A,f)$ obtained in (59), the second inequality from collapsing the geometric series, the third from the assumption on the scaling parameters contained in $\mathcal{L}$. Thus, combining the previous two estimates in (64) and (65) one has:
@@
$\leq Cd_{1}(x_{1},x^{\prime}_{1})^{2\alpha}d_{1}(x_{2},x^{\prime}_{2})^{2\alpha}\ldots d_{d}(x_{d},x^{\prime}_{d})^{2\alpha}$ (66)
Since the choice of $\mathbf{x},\mathbf{\tilde{x}}$ were arbitrarily chosen, $\Delta(A,f)\in\Lambda^{2\alpha}([0,1]^{d})$
Thus, from lemma 3, we can compute the approximation $\tilde{A}(f)=\sum_{\alpha=1}^{d}A^{\alpha}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\tilde{\mathbf{v}}^{\alpha}$ to $A(f)$ such that $\Delta(A,f)=A(f)-\tilde{A}(f)\in\Lambda^{2\alpha}([0,1]^{d})$.
3 Paraproduct Caricatures
3.1 3-tensor paraproducts and taylor expansion of $\tilde{A}(f)$
Corollary 1.
Suppose $A\in C^{3},f\in\Lambda^{2\alpha}([0,1]^{3})$, then we obtain the following discrete $3$-tensor paraproduct decomposition
$\tilde{A}(f)=\sum_{j_{1}=0_{1}}^{N_{1}}\sum_{j_{2}=0_{2}}^{N_{2}}\sum_{j_{3}=0_{3}}^{N_{3}}A^{\prime}(P^{j_{1},j_{2},j_{3}}(f))Q^{j_{1}}(f)Q^{j_{2}}(f)Q^{j_{3}}(f)+$
@@ 3 Paraproduct Caricatures
$+A^{\prime\prime\prime}(P^{j_{1},j_{2},j_{3}}(f))Q^{j_{3}}Q^{j_{2}}Q^{j_{1}}(f)$ (67)
@@
Corollary 2.
The approximation, $\tilde{A}(f)$, is real analytic on the image of $f$ and consequently can be expanded into a taylor series expansion for fixed scales $(j_{1},j_{2},\ldots,j_{d})$
$\tilde{A}(f)=\sum_{\alpha=1}^{d}A^{\alpha}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\tilde{\mathbf{v}}^{\alpha}$ (68)
where $A^{\alpha}(P^{j_{1},j_{2},\ldots,j_{d}}(f))$ is the derivative of A of order $\alpha$ evaluated at the average of $f$ and $\tilde{\mathbf{v}}^{\alpha}$ is an operator acting on $f$ containing wavelet coefficients in $\alpha$ directions.
@@
As any real analytic function can be represented by its taylor series, it suffices to show $A$ is real analytic on $f$. Since $\tilde{A}(f)=\sum_{\alpha=1}^{d}A^{\alpha}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\tilde{\mathbf{v}}^{\alpha}$ by Theorem 2 we argue $\sum_{\alpha=1}^{d}A^{\alpha}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\tilde{\mathbf{v}}^{\alpha}$ has a power series representation. $A^{\alpha}(P^{j_{1},j_{2},\ldots,j_{d}}(f))$ is a constant coefficient for $\alpha=1,2,\ldots,d$ and since $\tilde{\mathbf{v}}^{\alpha}$ represents $\alpha$ products of wavelet coefficients, $\sum_{\alpha=1}^{d}A^{\alpha}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\tilde{\mathbf{v}}^{\alpha}$ has a power series representation; therefore $A$ is real analytic on $f$ and has the taylor expansion $\sum_{\alpha=1}^{d}A^{\alpha}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\tilde{\mathbf{v}}^{\alpha}$ where $A^{\alpha}(P^{j_{1},j_{2},\ldots,j_{d}}(f))$ are the coefficients in the expansion which are derivatives of $A$. ∎
Corollary 3.
The sequence of operators $(\tilde{\mathbf{v}}^{1},\tilde{\mathbf{v}}^{2},\ldots,\tilde{\mathbf{v}}^{d})$ form a basis for the subspace of linear functionals of the form $U=\{g:g=A(f),f\in\Lambda^{\alpha}([0,1]^{d}),A\in C^{\infty}\}\subset V=\{h:h:[0,1]^{d}\to\mathbb{R}\}$ for a fixed sequences of scales $(j_{1},j_{2},\ldots,j_{d})$
Proof.
$V$ is a vectors space since for $h_{i},h_{j}\in V$ $h_{i}+h_{g}\in V$ and $\alpha h_{i}\in V$ for $\alpha\in\mathbb{R}$. $U\subset V$ is a subspace of $V$ by the same reasoning. Since $A(f)=\sum_{\alpha=1}^{d}A^{\alpha}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\tilde{\mathbf{v}}^{\alpha}$ by Theorem 2, the sequence of operators $(\tilde{\mathbf{v}}^{1},\tilde{\mathbf{v}}^{2},\ldots,\tilde{\mathbf{v}}^{d})$ span $U$. Suppose $A(f)=0$ such that $0=\sum_{\alpha=1}^{d}A^{\alpha}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\tilde{\mathbf{v}}^{\alpha}$. Then the operators $(\tilde{\mathbf{v}}^{1},\tilde{\mathbf{v}}^{2},\ldots,\tilde{\mathbf{v}}^{d})$ are linearly independent since $A^{1}(P^{j_{1},j_{2},\ldots,j_{d}}(f))=A^{2}(P^{j_{1},j_{2},\ldots,j_{d}}(f))=\cdots=A^{d}(P^{j_{1},j_{2},\ldots,j_{d}}(f))=0$ since the derivative of the 0 function is 0. Thus, since $(\tilde{\mathbf{v}}^{1},\tilde{\mathbf{v}}^{2},\ldots,\tilde{\mathbf{v}}^{d})$ span $U$ and are linearly independent, they form a basis for the subspace of linear functionals of the form: $U=\{g:g=A(f),f\in\Lambda^{\alpha}([0,1]^{d}),A\in C^{\infty}\}\subset V=\{h:h:[0,1]^{d}\to\mathbb{R}\}$
3.2 Calderon-Zygmund Decomposition
Theorem 3.
Consider the paraproduct decomposition $A(f)=\tilde{A}(f)+\Delta(A,f)$ as defined in Theorem 2. If we assume the average of $f$ with respect to each direction is 0, the paraproduct decomposition $A(f)=\tilde{A}(f)+\Delta(A,f)$, can be modified to obtain the following Calderon-Zygmund decomposition $A(f)=\tilde{A}(f)_{\lambda}+\Delta(A,f)_{\lambda}$, where $\tilde{A}(f)_{\lambda}$ is the bad part of $A(f)_{\lambda}$ and $\Delta(A,f)_{\lambda}$ is the good part of $A(f)$.
Proof.
In the interest of intelligibility, the analysis is carried out, without loss of generality, for a fixed sequence of scales $(j_{1},j_{2},\ldots,j_{d})$ and fixed order, $\alpha=m$, for the approximation term. First we provide a more explicit definition of the projection operator on the wavelet basis (from (18)) in terms of a series expansion of the wavelet coefficients supported on dyadic intervals, $k_{i}=1_{i},\ldots,2^{j_{i}}$, in direction i.
$Q^{j_{i}}(f):=\sum_{\tilde{k}_{i}=1_{i}}^{2^{j_{i}}+1}P_{\tilde{k}_{i}}^{{j_{i}}+1}(f)-\sum_{k_{i}=1_{i}}^{2^{j_{i}}}P_{k_{i}}^{j_{i}}(f)$
$=\sum_{k_{i}=1_{i}}^{2^{j_{i}}}Q^{j_{i}}_{k_{i}}(f)$ (69)
Similarly, a more explicit definition of $\tilde{\mathbf{v}}^{m}$, originally defined in (38), in terms of the wavelet coefficients defined on $d$-dimensional dyadic boxes is provided:
$\mathbf{v}_{k_{1},k_{2},\ldots,k_{d}}^{m}(f)\coloneq Q_{k_{d}}^{j_{d}}Q_{k_{d-1}}^{j_{d-1}}\ldots Q_{k_{m}}^{j_{m}}(f)Q_{k_{m-1}}^{j_{m-1}}(f)\ldots Q_{k_{1}}^{j_{1}}(f)$
$+Q_{k_{d}}^{j_{d}}Q_{k_{d-1}}^{j_{d-1}}\ldots Q_{k_{d-m+1}}^{j_{d-m+1}}(f)Q_{k_{d-m}}^{j_{d-m}}Q_{k_{m}}^{j_{m}}(f)\ldots Q_{k_{m-1}}^{j_{m-1}}(f)+\ldots+$
$Q_{k_{d}}^{j_{d}}(f)Q_{k_{d-1}}^{j_{d-1}}Q_{k_{d-2}}^{j_{d-2}}\ldots Q_{k_{d-m-1}}^{j_{d-m-1}}(f)Q_{k_{d-m-2}}^{j_{d-m-2}}(f)\ldots Q_{k_{1}}^{j_{1}}(f)$ (70)
$\tilde{\mathbf{v}}^{m}\coloneq\sum_{k_{1}=1_{1},k_{2}=1_{2},\ldots,k_{d}=1_{d}}^{2^{j_{1}},2^{j_{2}},\ldots,2^{j_{d}}}\mathbf{v}_{k_{1},k_{2},\ldots,k_{d}}^{m}(f)$
$=\sum_{k_{1}=1_{1},k_{2}=1_{2},\ldots,k_{d}=1_{d}}^{2^{j_{1}},2^{j_{2}},\ldots,2^{j_{d}}}Q_{k_{d}}^{j_{d}}Q_{k_{d-1}}^{j_{d-1}}\ldots Q_{k_{m}}^{j_{m}}(f)Q_{k_{m-1}}^{j_{m-1}}(f)\ldots Q_{k_{1}}^{j_{1}}(f)$
$+Q_{k_{d}}^{j_{d}}Q_{k_{d-1}}^{j_{d-1}}\ldots Q_{k_{d-m+1}}^{j_{d-m+1}}(f)Q_{k_{d-m}}^{j_{d-m}}Q_{k_{m}}^{j_{m}}(f)\ldots Q_{k_{m-1}}^{j_{m-1}}(f)+\ldots+$
$Q_{k_{d}}^{j_{d}}(f)Q_{k_{d-1}}^{j_{d-1}}Q_{k_{d-2}}^{j_{d-2}}\ldots Q_{k_{d-m-1}}^{j_{d-m-1}}(f)Q_{k_{d-m-2}}^{j_{d-m-2}}(f)\ldots Q_{k_{1}}^{j_{1}}(f)$ (71)
Its also profitable to define the set of all dyadic boxes at scales $(j_{1},j_{2},\ldots,j_{d})$:
$\mathcal{D}_{\lambda}\coloneq\{B_{k_{1},k_{2},\ldots,k_{d}}:B_{k_{1},k_{2},\ldots,k_{d}}\coloneq I^{j_{1}}_{k_{1}}\times I^{j_{2}}_{k_{2}}\times\ldots\times I^{j_{d}}_{k_{d}}\hskip 5.69046pt\forall\hskip 5.69046ptk_{1}=1_{1}\ldots 2^{j_{1}},k_{2}=1_{2}\ldots 2^{j_{2}},\ldots,$
$k_{d}=1_{d}\ldots 2^{j_{d}},B_{jk_{1},k_{2},\ldots,k_{d}}\coloneq\text{supp}(A^{m}(P_{k_{1},k_{2},\ldots,k_{d}}^{j_{1},j_{2},\ldots,j_{d}}(f))\mathbf{v}_{k_{1},k_{2},\ldots,k_{d}}^{m}(f)),$
$|A^{m}(P_{k_{1},k_{2},\ldots,k_{d}}^{j_{1},j_{2},\ldots,j_{d}}(f))\mathbf{v}_{k_{1},k_{2},\ldots,k_{d}}^{m}(f)|>\lambda\coloneq 2^{-(j_{1}+j_{2}+\ldots+j_{d})2\alpha}\}$ (72)
Let each $B_{k_{1},k_{2},\ldots,k_{d}}\in D_{\lambda}$ be indexed as $i=1,2,\ldots,|D_{\lambda}|$ such that each $B_{k_{1},k_{2},\ldots,k_{d}}\in D_{\lambda}$ corresponds to some $B_{i}\in D_{\lambda}$. Similarly, let $B_{k_{1},k_{2},\ldots,k_{d}}\in[0,1]^{d}\setminus D_{\lambda}$ be indexed by $p=1,2,\ldots,|B_{p}\in[0,1]^{d}\setminus D_{\lambda}|$ such that each $B_{k_{1},k_{2},\ldots,k_{d}}\in[0,1]^{d}\setminus D_{\lambda}$ corresponds to some $B_{p}\in[0,1]^{d}\setminus D_{\lambda}$. The set $D_{\lambda}$ contains all dyadic boxers where the $m$th term in the approximation larger than $\lambda$ for a fixed sequence of scales, $(j_{1},j_{2},\ldots,j_{d})$. We now define the purported Calderon-Zygmund decomposition in terms of the paraproduct decomposition by restricting the supports of the good and bad functions :
$A(f)=\tilde{A}(f)+\Delta(A,f)$
$=\sum_{B_{i}\in D_{\lambda}}\tilde{A}(f)+\Delta(A,f)+\sum_{B_{p}\in[0,1]^{d}\setminus D_{\lambda}}\tilde{A}(f)+\Delta(A,f)$
$=(\sum_{B_{i}\in D_{\lambda}}\tilde{A}_{i}(f))+(\sum_{B_{p}\in[0,1]^{d}\setminus D_{\lambda}}\Delta(A_{p},f))$
$=\tilde{A}(f)_{\lambda}+\Delta(A,f)_{\lambda}$ (73)
where we have defined the bad function as $\tilde{A}(f)_{\lambda}\coloneq(\sum_{B_{i}\in D_{\lambda}}\tilde{A}_{i}(f))$ and the good function as $\Delta(A,f)_{\lambda}\coloneq(\sum_{B_{p}\in[0,1]^{d}\setminus D_{\lambda}}\Delta(A_{p},f))$
From [9], if $A(f)\in L^{1}(\mathbb{R}^{d})$ and the following conditions are satisfied,
  1. 1.
    $A(f)=(\sum_{B_{i}\in D_{\lambda}}\tilde{A}_{i}(f))+\Delta(A,f)_{\lambda}$
  2. 2.
    $\lVert\Delta(A,f)_{\lambda}\rVert_{L^{\infty}([0,1]^{d})}\leq C\alpha,\lVert\Delta(A,f)_{\lambda}\rVert_{L^{1}([0,1]^{d})}\leq C\lVert A(f)\rVert_{L^{1}([0,1]^{d})}$
  3. 3.
    $\text{supp}(\tilde{A}_{i}(f)))\subset B_{i}$ , $\int\tilde{A}_{i}(f))=0\forall i=1,2,\ldots|D_{\lambda}|$, $\sum_{B_{i}\in D_{\lambda}}\lVert\tilde{A}_{i}(f)\rVert_{L^{1}([0,1]^{d})}\leq C\lVert A(f)\rVert_{L^{1}([0,1]^{d})}$
  4. 4.
    $\sum_{B_{i}\in D_{\lambda}}m(B_{i})\leq\frac{C}{\alpha}\lVert A(f)\rVert_{L^{1}}$
where $C>0,\alpha>0$, then the paraproduct decomposition $A(f)=\tilde{A}(f)+\Delta(A,f)$ can be represented as a Calderon-Zygmund decomposition:  $A(f)=\tilde{A}(f)_{\lambda}+\Delta(A,f)_{\lambda}$.
Since $A(f)$ is continuous with compact support, $A(f)\in L^{1}([0,1]^{d})$, satisfying the requisite condition on $A(f)$.Condition (1) follows immediately from Theorem, (2) and (73). The first part of condition (2) holds since:
$\lVert\Delta(A,f)_{\lambda}\rVert_{L^{\infty}([0,1]^{d})}=\lVert\displaystyle\sum_{B_{i}\in D_{\lambda}}\beta_{i}\chi_{B_{i}}(x_{1},x_{2},\ldots,x_{d})+\tilde{A}_{i}(f)\rVert_{L^{1}([0,1]^{d})}\leq C2^{-(j_{1}+j_{2}+\ldots+j_{d}))2\alpha}$ (74)
by the estimates on the coefficients, $\beta_{i}$, defined in (59) and by construction of $\tilde{A}_{i}(f)$. The second part of condition (2) holds since
$\lVert\Delta(A,f)_{\lambda}\rVert_{L^{1}([0,1]^{d})}=\int_{[0,1]^{d}}|(\sum_{B_{p}\in[0,1]^{d}\setminus D_{\lambda}}\Delta(A_{p},f))|dx$
$=\int_{[0,1]^{d}}|\sum_{B_{p}\in[0,1]^{d}\setminus D_{\lambda}}\tilde{A}(f)+\Delta(A,f)|dx$
$=\int_{B_{p}\in[0,1]^{d}\setminus D_{\lambda}}|\tilde{A}(f)+\Delta(A,f)|dx$
$\leq C\int_{[0,1]^{d}}|\tilde{A}(f)+\Delta(A,f)|dx$
$=\lVert A(f)\rVert_{L^{1}([0,1]^{d})}$ (75)
The first part of condition (3) by construction of the set $D_{\lambda}$ in (72), the second part of condition (3) by the assumption that the average of $f$ in each direction is 0, and finally the last part of condition 3 holds since:
$\displaystyle\sum_{B_{i}\in D_{\lambda}}\lVert\tilde{A}_{i}(f)\rVert_{L^{1}([0,1]^{d})}=\displaystyle\sum_{B_{i}\in D_{\lambda}}\int_{[0,1]^{d}}|\tilde{A}_{i}(f)|dx$
$=\displaystyle\sum_{B_{i}\in D_{\lambda}}\int_{B_{i}}|\tilde{A}_{i}(f)|dx$
$\leq C(\displaystyle\sum_{B_{i}\in D_{\lambda}}\int_{B_{i}}|\tilde{A}_{i}(f)|dx+\displaystyle\sum_{B_{p}\in[0,1]^{d}\setminus D_{\lambda}}\int_{B_{p}}|\tilde{A}_{i}(f)|)dx$
$=\lVert A(f)\rVert_{L^{1}([0,1]^{d})}$ (76)
where the last equality holds since $B_{i}\cap B_{p}=\emptyset$, and the inequality holds by the definition of $A(f)$. Finally condition (4) is met since:
$\sum_{B_{i}\in D_{\lambda}}m(B_{i})\leq C\int_{[0,1]^{d}}|A(f)|dx$
$=\lVert A(f)\rVert_{L^{1}([0,1]^{d})}$ (77)
which holds since $D_{\lambda}\subset[0,1]^{d}$, thus completing the proof.
4 Computational Examples
4.1 Quasilinearization of Heat and Potential Kernels (d=2)
As described in the previous section, the paraproduct and Calderon-Zygmund decompositions are connected; we illustrate this with tensor compression examples of the heat and potential kernels. The approximation captures the ”bad” part, (i.e. the part high frequency content) and has small support, while the residual captures the ”good” part (i.e. the low frequency content).
We process the original potential and heat kernels directly with the standard tensor haar basis, along with the paraproduct decomposition of both kernels. To do this, we randomly generated $\{x_{i}\}_{i=1}^{512}$ source and $\{y_{i}\}_{i=1}^{512}$ target points with each $y_{i},x_{i}\in[-1,1]^{3}$.We take $f(x,y)=\lVert x-y\rVert_{l_{2}}$. We organize $f$ to be $\alpha$-Holder using the questionnaire algorithm. We define $A_{1}(f)=\frac{1}{\lVert x-y\rVert_{l_{2}}}$ as the potential kernel and $A_{2}(f)=e^{-\frac{\lVert x-y\rVert_{l_{2}}}{\epsilon}}$ with $\epsilon=1$ as the heat kernel.
Figure 1: Points generated for heat and potential kernels
Figure 2: top row, left to right: $f,\tilde{A}_{2}(f),\Delta(A_{2},f)$ for heat kernel; bottom row, left to right: $f,\tilde{A_{1}}(f),\Delta(A_{1},f)$ for potential kernel
In Figure 2 we depict the original, approximation, and residual matrices of the heat and potential kernels. We process all of the matrices with the tensor Haar basis and plot the expansion coefficients in Figure 3 below.
Figure 3: (left) Heat kernel (right) Potential kernel expansion coefficients for original, approximation, and residual
In Figure 3 we plot the expansion coefficients for both the heat and potential kernels. In both cases, the matrices of the original functions, $A_{1}(f)$ and $A_{2}(f)$ exhibit slower decay than the approximation and residual matrices from the decomposition. This is quantified in the table below.
Kernels $|$ basis vectors 2000 6000 10000
Original heat kernel, $A_{1}(f)$ 36 % 72 % 91 %
Heat kernel approximation, $\tilde{A_{1}}(f)$ 80 % 98 % 99 %
Heat kernel residual $\Delta(A_{1},f)$ 66 % 91 % 97 %
Original potential kernel, $A_{2}(f)$ 40 % 75 % 92 %
Potential kernel approximation, $\tilde{A_{2}}(f)$ 80 % 98 % 99 %
Potential kernel residual, $\Delta(A_{2},f)$ 55 % 83 % 95 %
Table 1: Percentage of $l_{1}$ entropy captured by top $k\in\{{2000,6000,10000}\}$ basis vectors
Figure 3 and Table 1 demonstrate how the paraproduct decomposition connects to the Calderon-Zygmund decomposition; the high frequency content captured by the approximation has small support and can consequently be captured by only a few tensor basis vectors. The low frequency content captured by the residual has higher regularity ($2\alpha$-Holder vs $\alpha$-Holder) and consequently exhibits faster decay when processed by the tensor wavelet basis. This suggest downstream applications for construction of fast algorithms such as [6].
4.2 Cone singularity in complex plane (d=3)

### v2

@@ 3. Preliminaries
(59) $\leq Cd_{d}(x^{\prime}_{1},x_{1})^{2\alpha}d_{d}(x^{\prime}_{2},x_{2})^{2\alpha}\ldots d_{d}(x^{\prime}_{d},x_{d})^{2\alpha}$
since $\Delta_{(N_{i})_{i=1}^{d}}(A,f)$ is piecewise constant on dyadic boxes (see 54). Now consider the case where $x,x^{\prime}$ are not contained in same dyadic box. Again, summing over the scaling and location parameters in $\mathcal{L}$ for $\Delta_{(N_{i})_{i=1}^{d}}(A,f)$ one has:
@@
(60) $\leq Cd_{d}(x^{\prime}_{1},x_{1})^{2\alpha}d_{d}(x^{\prime}_{2},x_{2})^{2\alpha}\ldots d_{d}(x^{\prime}_{d},x_{d})^{2\alpha}$
where we order the scales in $\mathcal{L}$ such that $\mathbf{j}^{(0)}\leq\mathbf{j}^{(1)}\leq\ldots\leq\mathbf{j}^{(N)}$ indicates $2^{-(\sum_{l}^{d}\mathbf{j}_{l}^{(0)})}\leq 2^{-(\sum_{l}^{d}\mathbf{j}_{l}^{(1)})}\leq\ldots\leq 2^{-(\sum_{l}^{d}\mathbf{j}_{l}^{(N)})}$. The first inequality holds from the estimates on $\Delta_{(N_{i})_{i=1}^{d}}(A,f)$ obtained in (54), the second inequality from collapsing the geometric series, the third from the assumption on the scaling parameters contained in $\mathcal{L}$. Thus, combining the previous two estimates in (59) and (60) one has:
@@
(61) $\leq Cd_{d}(x^{\prime}_{1},x_{1})^{2\alpha}d_{d}(x^{\prime}_{2},x_{2})^{2\alpha}\ldots d_{d}(x^{\prime}_{d},x_{d})^{2\alpha}$
Since the choice of $\mathbf{x},\mathbf{\tilde{x}}$ were arbitrarily chosen, $\Delta_{(N_{i})_{i=1}^{d}}(A,f)\in\Lambda_{2\alpha}([0,1]^{d})$
Thus, from lemma 4.5, we can compute the approximation $\tilde{A}_{(N_{i})_{i=0}^{d}}(f)=\sum_{\beta=1}^{d}A^{\beta}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\tilde{\mathbf{v}}^{\beta}$ to $A(f)$ such that $\Delta_{(N_{i})_{i=1}^{d}}(A,f)=A(f)-\tilde{A}_{(N_{i})_{i=0}^{d}}(f)\in\Lambda_{2\alpha}([0,1]^{d})$.
5. Paraproduct Caricatures
Included in this section are caricatures of d-tensor paraproducts stated in the main theorem.
Corollary 5.1.
Suppose $A\in C^{3}(\mathbb{R}),f\in\Lambda_{2\alpha}([0,1]^{3})$, then we obtain the following discrete 33-tensor paraproduct decomposition
$\tilde{A}_{(N_{i})_{i=0}^{d}}(f)=\sum_{j_{1},j_{2},j_{3}=0}^{N_{1},N_{2},N_{3}}A^{\prime}(P^{j_{1},j_{2},j_{3}}(f))Q^{j_{1}}(f)Q^{j_{2}}(f)Q^{j_{3}}(f)+$
@@ 5. Paraproduct Caricatures
(62) $+A^{\prime\prime\prime}(P^{j_{1},j_{2},j_{3}}(f))Q^{j_{3}}Q^{j_{2}}Q^{j_{1}}(f)$
@@
Corollary 5.2.
The approximation, $\tilde{A}_{(N_{i})_{i=0}^{d}}(f)$, is real analytic on the image of $f$ and has the following taylor expansion
(63) $\tilde{A}_{(N_{i})_{i=0}^{d}}(f)=\sum_{\beta=1}^{d}A^{\beta}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\tilde{\mathbf{v}}^{\beta}$
where $A^{\beta}(P^{j_{1},j_{2},\ldots,j_{d}}(f))$ are the derivatives of A of order $\beta$ evaluated at the average of $f$ and $\tilde{\mathbf{v}}^{\beta}$ are wavelet operators acting in $\beta$ directions.
@@
As any real analytic function can be represented by its taylor series, it suffices to show AA is real analytic on $f$. Since $\tilde{A}_{(N_{i})_{i=0}^{d}}(f)=\sum_{\beta=1}^{d}A^{\beta}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\tilde{\mathbf{v}}^{\beta}$ by Theorem 4.1 the power series representation $\sum_{\beta=1}^{d}A^{\beta}(P^{j_{1},j_{2},\ldots,j_{d}}(f))\tilde{\mathbf{v}}^{\beta}$ is immediate, as $\beta$ represents the derivatives of $A$ and the powers of $\tilde{\mathbf{v}}$
6. Cone singularity in complex plane (d=3)